The sum of Tray's age and Ja’s age is 60. Nine years ago, Tray was twice as old as Ja then. How old is Ja?
Answer by ikleyn(53763)

.

            It is simpler to solve this problem  MENTALLY,  without using equations,  than with equations.


Nine years ago, the sum of their ages was  60-2*9 = 42 years,

and Tray was twice the Ja's age.


Then it is OBVIOUS that the Ja's   age was  42/3 = 14 years then, and

                        the Tray's age was  14*2 = 28 years.


Hence, now they are 14+9 = 23 years old  (Ja's age)  and  28+9 = 37 years old (Tray's age).    ANSWER

Solved.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.